# Exact mobility edges in quasiperiodic network models with slowly varying potentials

### v1

@@ II.3 Analytical solutions from effective Hamiltonian
Substituting these solutions into the equation for $u_{i}$ will yields
@@
This equation will involves only sites $u_{i}$ and $u_{i\pm 2}$, thus the periodic sites of $u_{i\pm 1}$ are integrated out. In this way, we can obtain an equivalent equation involving only quasiperiodic sites. Furthermore, let us re-define the operators using $u_{i}\rightarrow u_{j}$, with $j=[i/2]$, assuming that $i\in 2\mathbb{Z}$, and have
@@
$g(E)V_{j}u_{j}+t^{2}u_{j-1}+t^{2}u_{j+1}=f(E)u_{j}.$ (20)
where $f(E)=E^{2}-2t^{2}$ and $g(E)=E$. This equation is the same as the original quasiperiodic slowly varying model, however, the corresponding coefficients have been changed. It has been well-known that the MEs of this model happens at $E_{c}=\pm(2t-V)$, with energy spectra boundaries at $E_{b}=\pm(2t+V)$, we can immediately find that in our model the MEs and spectra boundaries for Eq. (20) happen at
$f(E)=\pm(2t\pm g(E)V).$ (21)
Substituting $f(E)$ and $g(E)$ into the above equation will yield eight solutions, four of which are MEs given by Eq. (5), two of them are resonant states with $E=0$, and the other two for spectra boundaries at
$E_{b}=\pm(\sqrt{V^{2}+16t^{2}}+V)/2.$ (22)
Similarly, for any $\kappa$, or considering various nonzero potentials regardless of $V_{i}=0$ at the periodic sites, the system can always be mapped to an effective model like Eq. (20) with various $f(E)$ and $g(E)$, and one of the central conclusions in this work is that its MEs are located at Eq. (21). We will show below that when $V\rightarrow\infty$, $E_{b}=\pm V$, which has clear physical interpretation since the wave functions are localized at the extrema of the potential.
II.4 Analytical solutions from semianalytical method
Then, we understand the above results using the semi-analytical method. Starting with the effective potential obtained in the previous subsection, we follow the method outlined by Das Sarma et al. [49]. In the following, we consider the case with $\kappa=2$, and the similar analysis can be extended to the other models with slowing varying potentials.
Figure 6: Function $\cos(\pi\alpha j^{\nu})$ with (a) $\nu=1.6$ and (b) $\nu=0.6$. (c) and (d) show $s=(E-c)E$ with $c>0$ and $c<0$, respectively, and their corresponding regime for $40<s<$ 4. The red dots denote the MEs.
We begin with the eigenvalue equations
$-tu_{i-2}-tu_{i}$ $=$ $Eu_{i-1},$ (23)
$-tu_{i-1}+V_{i}u_{i}-tu_{i+1}$ $=$ $Eu_{i},$ (24)
$-tu_{i}-tu_{i+2}$ $=$ $Eu_{i+1},$ (25)
for the three sites from $(i-1)$ to $(i+1)$. Integrating out the sites with zero potential leads to the effective eigenvalue equation as
$t^{2}u_{j-1}+\lambda_{j}u_{j}+t^{2}u_{j+1}=f(E),$ (26)
with effective potential and effective eigenenergy
$\lambda_{j}=EV{\rm cos}(2^{\nu}\pi\alpha j^{\nu}+\phi),\quad f(E)=E^{2}-2t^{2}.$ (27)
The effective potential $\lambda_{j}$ exhibits the property of becoming locally constant in the thermodynamic limit
$\lim\limits_{j\rightarrow\infty}|\frac{d\lambda_{j}}{dj}|=\lim\limits_{j\rightarrow\infty}-EV2^{\nu}\pi\alpha\nu\frac{{\rm sin}(2^{\nu}\pi\alpha j^{\nu}+\phi)}{j^{1-\nu}}=0,$ (28)
due to $10<\nu<$ 1 [see Fig. 6(b)]. When $\nu>1$, the oscillating of the potential will become more rapid, yielding fully localized states in the thermodynamic limit [see Fig. 6(a)].
For the eigenstates of slowly varying models, one always expect its wave function with amplitude
@@ II.4 Analytical solutions from semianalytical method
$V_{j}=V{\rm cos}(2^{\nu}\pi\alpha j^{\nu}+\phi),$ (31)
which satisfies $|V_{j}|\leq V$. The crucial point is that when $V_{j}$ is locally constant for large $j$, in the vicinity of this regime we can treat it to be a constant, that is,
@@
with $-V\leq c\leq V$, yielding the following simplified solution
$t^{2}z^{2}+[(E-c)E-2t^{2}]z+t^{2}=0,,$ (33)
@@
Based on this equation, Das Sarma et al. [49] pointed out that the real-complex transition points of the above eigenvalue equation are the extended-localized transition points of the quasiperiodic model. If $\Delta<0$, the amplitude $z$ is complex, satisfying $|z|=1$, which means that the state is extended. Then we can conclude that the conditions for extended states are
@@
For $E>0$, the conditions become
$[(E-c)E]_{\text{max}}$ $=[(E+V)E]<4t^{2},$
$[(E-c)E]_{\text{min}}=[(E$ $-V)E]>0\Rightarrow E>V,$ (37)
while for $E<0$, they become
$[(E-c)E]_{\text{max}}$ $=[(E-V)E]<4t^{2},$
$[(E-c)E]_{\text{min}}=[(E$ $+V)E]>0\Rightarrow E<-V.$ (38)
Combining Eqs. (37) and (38), the exact MEs are determined by Eq. (5) [see Figs. 6(c) and 6(d)].
Figure 7: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for multipath quasiperiodic chain with slowly varying potentials when (a) $\gamma=2$, (b) $\gamma=3$, (c) $\gamma=4$, and (d) $\gamma=5$. The solid lines denote the MEs described by $f(E)=\pm(2t\pm g(E)V)$ with function $f(E)$ and $g(E)$ given in Eq. (43). The other parameters are the same with Fig. 2.
@@ III.1 Hermitian models
The method of solving the MEs in quasiperiodic mosaic models with slowly varying potentials through effective Hamiltonian can be generalized to various quasiperiodic network models with slowly varying potentials [35]. Consider a concrete case that can be described by the Hamiltonian
$\mathcal{H}$ $=\sum_{i=1}^{L}(V_{i}c_{i}^{\dagger}c_{i}+\sum_{k=1}^{\gamma}U_{k}d_{i,k}^{\dagger}d_{i,k})-t(\sum_{i=1}^{L}\sum_{k=1}^{\gamma}c_{i}^{\dagger}d_{i,k}$
$+\sum_{i=1}^{L-1}\sum_{k=1}^{\gamma}c_{i+1}^{\dagger}d_{i,k}+\mathrm{H.c.}),$ (39)
where $d_{i,k}^{\dagger}$ ($d_{i,k}$) is the creation (annihilation) operator at the $k$-th site of cell $i$, $V_{i}$ denotes the quasiperiodic slowly varying potentials defined as $V_{i}=V\cos(\pi\alpha i^{\nu})$, and $U_{k}$ represents constant potentials. Here we set $U_{k}=0$ for simplicity, and the results for nonzero cases are similar. Figure 2(b) illustrates the multipath quasiperiodic network model with $\gamma=3$. For such a multipath quasiperiodic model, the MEs can also be obtained through the effective Hamiltonian. To see that, we begin with the eigenvalue equations for the periodic sites
@@
where |$\ket{\psi^{n}}=\sum_{i}(u_{i}^{n}c_{i}^{\dagger}+\sum_{k}v_{i,k}^{n}d_{i,k}^{\dagger})\ket{0}$. Substituting these into the eigenvalue equation for $u_{i}$ yields
$(E-V_{i})u_{i}=\frac{t^{2}\gamma(u_{i-1}+2u_{i}+u_{i+1})}{E},$ (41)
Rewrite $u_{i}$ as $u_{j}$ to distinguish between the effective chain and original chain, we obtain the effective Hamiltonian
$\sum_{j}g(E)V\cos(\pi\alpha j^{\nu})u_{j}+t^{2}u_{j-1}+t^{2}u_{j+1}=f(E)u_{j},$ (42)
@@
$g(E)=\frac{E}{\gamma},\hskip 14.22636ptf(E)=-2t^{2}+\frac{E^{2}}{\gamma},$ (43)
which leads to the MEs at $f(E)=\pm(2t\pm g(E)V)$.
In Fig. 7, the fractal dimension of the eigenstates is plotted as a function of energy and potential strength for various $\gamma$. The black lines represent the MEs described by $f(E)=\pm(2t\pm g(E)V)$, with the corresponding $f(E)$ and $g(E)$ given in Eq. (43). It is observed that the value of $\gamma$ doesn’t affect the number of MEs, which remains fixed at $4$ for various $\gamma$, in contrast to the influence of $\kappa$. Additionally, there always exists a state at $E=0$, i.e., resonant state, which remains extended regardless of $V$.
Figure 8: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for PT-symmetric multipath quasiperiodic chain with slowly varying potentials when (a) $\gamma=1$, $\Gamma=0.5t$, (b) $\gamma=1$, $\Gamma=2t$, (c) $\gamma=2$, $\Gamma=0.5t$, and (d) $\gamma=2$, $\Gamma=2t$. The solid lines denote the MEs described by $f(E)=\pm(2t\pm g(E)V)$ with function $f(E)$ and $g(E)$ given in Eq. (48). The other parameters are the same with Fig. 2.
@@ III.2 Non-Hermitian models
$\mathcal{H}$ $=\sum_{i=1}^{L}(V_{i}c_{i}^{\dagger}c_{i}+\sum_{k=1}^{\gamma}i\Gamma d_{i,2k}^{\dagger}d_{i,2k}-i\Gamma d_{i,2k-1}^{\dagger}d_{i,2k-1})$
$-t(\sum_{i=1}^{L}\sum_{k^{\prime}=1}^{2\gamma}c_{i}^{\dagger}d_{i,k^{\prime}}+\sum_{i=1}^{L-1}\sum_{k^{\prime}=1}^{2\gamma}c_{i+1}^{\dagger}d_{i,k^{\prime}}+\mathrm{H.c.}).$ (44)
The PT symmetry is realized by balanced gain and loss in the periodic sites. The eigenvalue equations for the periodic sites are
@@
Substituting these equations into the eigenvalue equation for $u_{i}$, we obtain
$(E-V_{i})u_{i}=\frac{2\gamma Et^{2}(u_{i-1}+2u_{i}+u_{i+1})}{E^{2}+\Gamma^{2}}.$ (46)
Figure 9: (a) An illustration of the experimental setup based on a waveguide system for quasiperiodic mosaic model with $\kappa=2$. Here, the blue waveguides have quasiperiodic slowly varying potentials, while the red ones have zero potential. (b) Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for $L=100$. The solid lines denote the MEs described by Eq. (5). Evolution of the wave function $\psi(i,z)$ for the initial state $\psi(i,0)=\delta_{i,50}$ when (c) $V=0.5t$, (d) $V=t$, (e) $V=2t$, and (f) $V=3t$.
Rewrite $u_{i}$ using $u_{j}$ leads to the following effective Hamiltonian
$\sum_{j}g(E)V\cos(\pi\alpha j^{\nu})u_{j}+t^{2}u_{j-1}+t^{2}u_{j+1}=f(E)u_{j},$ (47)
@@
$g(E)=\frac{(E^{2}+\Gamma^{2})}{2\gamma E},\hskip 14.22636ptf(E)=-2t^{2}+\frac{E^{2}+\Gamma^{2}}{2\gamma}.$ (48)
This solution will yields MEs at $f(E)=\pm(2t\pm g(E)V)$, which is numerically verified in Fig. 8 for various $\gamma$ and $\Gamma$. The results for the other non-Hermitian cases can be derived in the same way.
@@ IV Experimental Proposal
The realization of quasiperiodic network models with various periodic network sites is challenging in cold-atom systems [52, 53, 54], thus we propose an experimental scheme based on coupled optical single-mode waveguides [55, 56, 17, 57]. By controlling the distance between neighboring waveguides and tuning their refraction index, we can modulate the hopping strength and on-site potential [58, 59], enabling the realization of various quasiperiodic network models with slowly varying potentials. To observe the localized-delocalized transition in such systems, we can inject light into a single waveguide and measure the distribution of light intensity after it has propagated through a certain distance in the waveguides array [see Fig. 9(a)]. The propagation of light along the waveguide can be described by tight-binding equation
@@
In Fig. 9(b), we plot the fractal dimension of the eigenstates as a function of energy $E$ and potentail strength $V$ for a small system ($L=100$), where we can also see clear MEs. Here we consider quasiperiodic mosaic model with slowly varying potentials when $\kappa=2$ as an example. To simulate the experimental results, we plot the evolution of the wavefunction for various potential strength $V$ in Figs. 9(c)-9(f). Although the specific wavefunction evolution depends on the initial state, the properties of localized state or extended state remain the same. Here, we set the initial state as $\psi(i,0)=\ket{0,\cdots,1_{i=50},\cdots,0}$. For $V=0.5t$, the wave function gradually spreads through the entire system. As the potential strength increases to $V=t$, the wavefunction propagating to the boundary decreases, since the number of extended states in the system is reduced. When $V=2t$ and $V=3t$, the wavefunctions become localized around the center of the system, as all states away from the band center are localized. These results provide some theoretical basis for future experiments.
Here, we only demonstrate the possible observation of localized-delocalized transition with the increasing of potential strength in network models in some small systems. The precise phase boundary from this kind of measurement can’t be precisely determined in the current stage, which is a common problem in all quantum simulating systems [60, 61, 62]; see our discussion in the next section. New algorithm is required to address this difficulty, which will be explored in future.

### v2

@@ II.3 Analytical solutions
Substituting these expressions into the equation for $u_{i}$ yields
@@
This equation involves only sites $u_{i}$ and $u_{i\pm 2}$, thus the periodic sites of $u_{i\pm 1}$ are integrated out. In this way, we can obtain an equivalent equation involving only quasiperiodic sites. Furthermore, let us re-define the operators using $u_{i}\rightarrow u_{j}$, with $j=[i/2]$, assuming that $i\in 2\mathbb{Z}$, we have
@@
$\lambda_{j}u_{j}+t^{2}u_{j-1}+t^{2}u_{j+1}=f(E)u_{j}.$ (23)
where
$\lambda_{j}=g(E)V_{j},\quad g(E)=E,\quad f(E)=E^{2}-2t^{2}.$ (24)
With effective potential strength $g(E)V$ and eigenenergy $f(E)$, we can directly obtian the MEs.
Figure 6: Function $\cos(\pi\alpha j^{\nu})$ with (a) $\nu=1.6$ and (b) $\nu=0.6$. (c) Function $s=(E-c)E$ with $c>0$ (left panel) and $c<0$ (right panel), and their corresponding regime for $0<s<4t^{2}$, with $t=1$. The red dots denote the MEs.
Next we provide a more detailed explanation of the above results using the method developed by Das Sarma et al. [59]. The effective potential $\lambda_{j}$ exhibits the property of becoming locally constant in the thermodynamic limit
$\lim\limits_{j\rightarrow\infty}|\frac{d\lambda_{j}}{dj}|=\lim\limits_{j\rightarrow\infty}-EV\pi\alpha\nu\frac{{\rm sin}(\pi\alpha j^{\nu}+\phi)}{j^{1-\nu}}=0,$ (25)
due to $10<\nu<$ 1 [see Fig. 6(b)]. In contrast when $\nu>1$, the oscillating of the potential will become more and more rapid, yielding fully localized states in the thermodynamic limit [see Fig. 6(a)].
For the eigenstates of slowly varying models, one always expects the wave function with amplitude
@@
$V_{j}=V{\rm cos}(\pi\alpha j^{\nu}+\phi),$ (28)
which satisfies $|V_{j}|\leq V$. The crucial point is that when $V_{j}$ is locally constant for large $j$, in the vicinity of this regime, we can treat it to be a constant, that is
@@
with $-V\leq c\leq V$, yielding the following simplified equation
$t^{2}z^{2}+[(E-c)E-2t^{2}]z+t^{2}=0,$ (30)
@@
Based on this equation, Das Sarma et al. [59] pointed out that the real-complex transition of the above eigenvalue determine the MEs for extended-localized transition. Obviously, when $\Delta<0$, the amplitude $z$ is complex, satisfying $|z|=1$, which means that the state is extended; otherwise, it is a real value with $|z|\neq 1$, which gives the localized state. Then we can find that the condition for extended states at
@@
Obviously, it is a quadratic function of $E$, yielding the curve in Fig. 6(c). When $c=V$, we find
$E={1\over 2}(V\pm\sqrt{16t^{2}+V^{2}}),\quad E=0,\quad E=V,$ (34)
and when $c=-V$, we have
$E={1\over 2}(-V\pm\sqrt{16t^{2}+V^{2}}),\quad E=0,\quad E=-V.$ (35)
These solutions correspond the MEs given by Eq. (6), resonant states $E=0$, and energy spectrum boundaries.
The above results have an immediate corollary, that is, the phase boundaries are independent of the phase carried in the potential, assuming $V_{j}=V\cos(\pi\alpha j^{\nu}+\phi)$. The reason is that while this phase will influence the potential at each site, it will not influence the maximal and minimal value of the potential. Meanwhile, the parameters of $\alpha$ is also not important, nor it is not necessary to be an irrational number. We find that $\alpha$ and $\phi$ will never appear in the expression of MEs. These are also verified by our numerical simulation, which are not shown.
Figure 7: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for multipath chain with slowly varying potentials when (a) $N=2$, (b) $N=3$, (c) $N=4$, and (d) $N=5$. The solid lines denote the MEs described by $f(E)=\pm(2t^{2}\pm g(E)V)$ with function $f(E)$ and $g(E)$ given in Eq. (40). The other parameters are the same as Fig. 2.
@@ III.1 Hermitian models
The above method with effective Hamiltonian can be used to determine the MEs in more generalized quasiperiodic network models with slowly varying potentials [28]. Consider a concrete case that can be described by the Hamiltonian
$\mathcal{H}$ $=\sum_{i=1}^{L}(V_{i}c_{i}^{\dagger}c_{i}+\sum_{k=1}^{N}U_{k}d_{i,k}^{\dagger}d_{i,k})-t(\sum_{i=1}^{L}\sum_{k=1}^{N}c_{i}^{\dagger}d_{i,k}$
$+\sum_{i=1}^{L-1}\sum_{k=1}^{N}c_{i+1}^{\dagger}d_{i,k}+\mathrm{H.c.}),$ (36)
where $d_{i,k}^{\dagger}$ ($d_{i,k}$) is the creation (annihilation) operator at the $k$-th site of cell $i$, $V_{i}$ denotes the quasiperiodic slowly varying potentials defined as $V_{i}=V\cos(\pi\alpha i^{\nu})$, and $U_{k}$ represents constant potentials. Here we set $U_{k}=0$ for simplicity, and the results for nonzero cases are similar. Fig. 2(b) illustrates the multipath quasiperiodic network model with $N=3$. For such a multipath quasiperiodic model, the MEs can also be obtained through the effective Hamiltonian. To see that, we begin with the eigenvalue equations for the periodic sites
@@
where |$\ket{\psi^{n}}=\sum_{i}(u_{i}^{n}c_{i}^{\dagger}+\sum_{k}v_{i,k}^{n}d_{i,k}^{\dagger})\ket{0}$. Substituting these results into the eigenvalue equation for $u_{i}$ yields
$(E-V_{i})u_{i}=\frac{t^{2}N(u_{i-1}+2u_{i}+u_{i+1})}{E}.$ (38)
Further, let us rewrite $u_{i}$ as $u_{j}$ to distinguish the index between the effective chain and original chain, we can obtain the effective Hamiltonian
$\sum_{j}g(E)V_{j}u_{j}+t^{2}u_{j-1}+t^{2}u_{j+1}=f(E)u_{j},$ (39)
@@
$g(E)=\frac{E}{N},\hskip 14.22636ptf(E)=-2t^{2}+\frac{E^{2}}{N},$ (40)
which leads to the MEs at $f(E)=\pm[2t^{2}\pm g(E)V]$.
Next, we verify the above results using numerical method. In Fig. 7, the fractal dimension $D_{2}$ of the eigenstates is plotted as a function of energy and potential strength for various $N$. The black lines represent the MEs described by $f(E)=\pm(2t^{2}\pm g(E)V)$, with the corresponding $f(E)$ and $g(E)$ given by Eq. (40). This equation have four solutions, thus four MEs, which is independent of $N$. Additionally, there always exists a state at $E=0$ as a resonant state, which remains extended regardless of $V$. This state is given by $g(E)=0$.
Figure 8: Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for PT-symmetric multipath chain with slowly varying potentials when (a) $N=1$, $\Gamma=0.5t$, (b) $N=1$, $\Gamma=2t$, (c) $N=2$, $\Gamma=0.5t$, and (d) $N=2$, $\Gamma=2t$. The solid lines denote the MEs described by $f(E)=\pm(2t^{2}\pm g(E)V)$ with function $f(E)$ and $g(E)$ given in Eq. (45). The other parameters are the same as Fig. 2.
@@ III.2 Non-Hermitian models
$\mathcal{H}$ $=\sum_{i=1}^{L}(V_{i}c_{i}^{\dagger}c_{i}+\sum_{k=1}^{N}i\Gamma d_{i,2k}^{\dagger}d_{i,2k}-i\Gamma d_{i,2k-1}^{\dagger}d_{i,2k-1})$
$-t(\sum_{i=1}^{L}\sum_{k^{\prime}=1}^{2N}c_{i}^{\dagger}d_{i,k^{\prime}}+\sum_{i=1}^{L-1}\sum_{k^{\prime}=1}^{2N}c_{i+1}^{\dagger}d_{i,k^{\prime}}+\mathrm{H.c.}).$ (41)
The PT symmetry is realized by balancing the gain and loss in the periodic sites. The eigenvalue equations for the periodic sites are
@@
Substituting these results into the eigenvalue equation of $u_{i}$ will yield
$(E-V_{i})u_{i}=\frac{2NEt^{2}(u_{i-1}+2u_{i}+u_{i+1})}{E^{2}+\Gamma^{2}}.$ (43)
Figure 9: (a) An illustration of the experimental setup based on optical waveguides for quasiperiodic mosaic model with $\kappa=2$. Here, the blue waveguides represent the quasiperiodic sites with slowly varying potentials, while the red ones represent the periodic sites with zero potential. (b) Fractal dimension $D_{2}$ of different eigenstates as a function of energy $E$ and potential strength $V$ for $L=100$. The solid lines denote the MEs described by Eq. (6). Evolution of the wave function $|\psi(i,z)|$ for the initial state $\psi(i,0)=\delta_{i,50}$ when (c) $V=0.5t$, (d) $V=t$, (e) $V=2t$, and (f) $V=3t$.
Similarly, let us rewrite $u_{i}$ using $u_{j}$ and we can obtain the following effective Hamiltonian
$\sum_{j}g(E)V_{j}u_{j}+t^{2}u_{j-1}+t^{2}u_{j+1}=f(E)u_{j},$ (44)
@@
$g(E)=\frac{(E^{2}+\Gamma^{2})}{2NE},\hskip 14.22636ptf(E)=-2t^{2}+\frac{E^{2}+\Gamma^{2}}{2N}.$ (45)
This solution will yields MEs at $f(E)=\pm(2t^{2}\pm g(E)V)$, which is numerically verified for various $N$ and $\Gamma$ in Fig. 8. The results for the other non-Hermitian cases can be derived in the same way.
@@ IV Experimental Proposal
Finally, we discuss the possible way to realize these network models. The realization of quasiperiodic network models with various periodic network sites is challenging in cold-atom systems [64, 65, 66], however, it can be easily realized using the coupled single-mode optical waveguides [67, 68, 19, 69]. By controlling the distance between neighboring waveguides and tuning their refraction index, we can modulate the hopping strength and on-site potential [70, 71], enabling the realization of various quasiperiodic network models with slowly varying potentials. To observe the localized-delocalized transition in such systems, we can inject light into a single waveguide and measure the distribution of light intensity after it propagates through a certain distance in the waveguides array [see Fig. 9(a)]. The propagation of light along the waveguide can be described by tight-binding equation
@@
In Fig. 9(b), we plot the fractal dimension of the eigenstates as a function of energy $E$ and potentail strength $V$ for a small system ($L=100$), where we can also see clear MEs. Here we consider quasiperiodic mosaic model with slowly varying potentials when $\kappa=2$. To simulate the experimental results, we plot the evolution of the wave function for various potential strength $V$ in Figs. 9(c) - 9(f). Although the specific evolution of wave function depends strongly on its initial state, the properties of localized state or extended state remain the same. Here, we set the initial state as $\psi(i,0)=\ket{0,\cdots,1_{i=50},\cdots,0}$. For $V=0.5t$, the wave function gradually spreads through the entire system. As the potential strength increases to $V=t$, the wave function propagating to the boundary decreases, since the number of extended states in the system is reduced. When $V=2t$ and $V=3t$, the wave functions become localized around the center of the system, as all states away from the band center are localized, which we leave as an open question to be investigated in the future. Our recent experiments about the reentrant phenomena in quasiperiodic systems can be find in Ref. [63].
Here, we only demonstrate the possible observation of localization-delocalization transition with the increasing of potential strength in network models in some small systems. The precise phase boundary from this kind of measurement can not be precisely determined in the current stage, which is a common problem in all quantum simulating systems [72, 73, 74]; see our discussion in the next section. To this end, new algorithm to exact the MEs from the evolution is required in future.